High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor-hours. At the beginning of the year, the company provided the following estimates: Department Molding Painting Direct labor-hours 12,000 60,000 Machine-hours 70,000 8,000 Fixed manufacturing overhead cost $ 497,000 $ 615,000 Variable manufacturing overhead per machine-hour $ 1.50 - Variable manufacturing overhead per direct labor-hour - $ 2.00 Job 205 was started on August 1 and completed on August 10. The company's cost records show the following information concerning the job: Department Molding Painting Direct labor-hours 30 84 Machine-hours 110 20 Direct materials $ 770 $ 1,332 Direct labor cost $ 525 $ 1,470 Required: 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. Compute the total overhead cost applied to Job 205. 3-a. What would be the total manufacturing cost recorded for Job 205? 3-b. If the job contained 50 units, what would be the unit product cost?
Answer #1

1) Predetermine overhead rate :

Molding department rate = (497000/70000+1.50) = 8.60 per MH

Painting department rate = (615000/60000+2) = 12.25 per DLH

2) Total overhead cost applied = (110*8.60+84*12.25) = 1975

3a) Total manufacturing cost

Direct material (770+1332) 2102
Direct labor (525+1470) 1995
Overhead 1975
Total manufacturing cost 6072

3b) Unit product cost = 6072/50 = 121.44